OpenStudy (anonymous):

You are tiling a kitchen floor that is 10 ft wide by 4 yd long. How many square yards of tile do you need?

OpenStudy (helder_edwin):

convert feet to yards or yards to feet. then multiply them

OpenStudy (anonymous):

idk how to do that..

OpenStudy (helder_edwin):

1 yd = 3 ft

OpenStudy (helder_edwin):

so 4 yd = 4*3 ft=12 ft

OpenStudy (helder_edwin):

so u need 10 ft * 12 ft=120 ft^2 of tiles

OpenStudy (helder_edwin):

1 yd = 3 ft then 1ft=1/3 yd

OpenStudy (anonymous):

the answer choices are: A) 13 1/3 yd^2 B) 13 1/2 yd^2 C)20 yd^2 D) 40 yd^2

OpenStudy (helder_edwin):

so 120 ft^2=120 (1/3 yd)^2=120/9 yd^2

OpenStudy (anonymous):

wut..

OpenStudy (anonymous):

that makes no sence to me..

OpenStudy (helder_edwin):

120/9 yd^2=13 1/3 yd

OpenStudy (helder_edwin):

13 1/3 yd^2

OpenStudy (anonymous):

ooohhh haha thank youu! (:

OpenStudy (helder_edwin):

u r welcome